# X-Reflect: Cross-Reflection Prompting for Multimodal Recommendation

### v1

@@ Results
MovieLens-1M Amazon-Software Precision@10 Recall@10 NDCG@10 Precision@10 Recall@10 NDCG@10 X-Reflect 0.3111 0.2746 0.3900 0.0431 0.3969 0.3024 X-Reflect-keyword 0.3128 0.2754 0.3908 0.0425 0.3917 0.3125
Discussions and Conclusions
In this paper, we introduce a novel framework, X-Reflect, which leverages cross-reflection prompting to enrich item descriptions by integrating both textual and visual modalities. Through a series of comprehensive experiments on two widely recognized recommendation benchmarks–MovieLens-1M and Amazon-Software–we demonstrate that our approach consistently outperforms several baseline prompting strategies, particularly those that do not incorporate cross-reflection.
Our findings suggest that by prompting Large Multimodal Models (LMMs) to explicitly identify and reconcile supportive or conflicting information between text and image, X-Reflect is able to generate more diverse and contextually rich item representations. This, in turn, enhances the effectiveness of recommendation systems by providing a more holistic understanding of the items being recommended.
@@
However, the proposed framework is not without its limitations. One notable drawback is the increased computational overhead associated with the need for multiple LMM inferences per item. This requirement can be resource-intensive, especially when processing large-scale datasets or deploying the system in real-time applications. To mitigate this issue, future research could explore optimization strategies, such as selectively applying cross-reflection prompting to a subset of items that are likely to benefit the most from this process. Additionally, advances in model efficiency, such as improved inference algorithms or more lightweight LMM architectures, could further reduce the computational burden.
Despite these challenges, the potential of cross-reflection prompting to enhance multimodal recommendations is significant. By continuing to refine and optimize this approach, we believe it can be a valuable tool in the ongoing effort to develop more accurate, personalized, and context-aware recommendation systems. Future work will also explore the integration of supervised fine-tuning and continuous pre-training to further improve the quality of responses generated by LMMs, as well as investigating the types of augmented information that are most beneficial for various downstream tasks.
Ethical Considerations
Recommendation systems, particularly those utilizing multimodal prompting methods like ours, significantly influence user behavior. There is a risk that such systems could inadvertently promote biased or harmful content or products, especially if the context of both visual and textual data is not properly assessed. For instance, our method, which integrates information from both images and text, might recommend items that are visually appealing but are contextually inappropriate or unsafe if the nuances of the data are misinterpreted.
To mitigate these risks, it is crucial to incorporate robust content filtering mechanisms within the multimodal framework to detect and block the promotion of harmful or inappropriate content. This includes not only evaluating the textual data for potential issues but also ensuring that the visual data aligns with safety standards. Additionally, our system should be designed with a strong emphasis on user well-being, avoiding recommendations that could lead to negative outcomes, such as encouraging excessive spending, unhealthy behaviors, or exposure to harmful content. By prioritizing safety and ethical considerations in our multimodal prompting approach, we can better protect users and promote positive interactions with the recommendation system.
Appendix A Appendix
KAR Augmentation Details
In the work by Xi et al. (2023), a specific prompt was designed to elicit factual knowledge about items. The prompt instructed the model to: “Introduce {item}, {item description} and describe its attributes precisely (including but not limited to {scenario-specific factors})”. The {scenario-specific factors} were identified by prompting an LLM with the following instruction: “List the importance factors or features that determine whether a user will be interested in a {item}.”
For the MovieLens-1M dataset, we directly use the factual knowledge about movies generated in their study. For the Amazon-Software dataset, we apply the same approach to first identify the {scenario-specific factors}. The prompt was adapted to: “List the importance factors or features that determine whether a user will be interested in software.” The set of factors generated by ChatGPT includes functionality, user interface, user experience, performance, compatibility, cost, security, support and documentation, customizability, scalability, reliability, updates and maintenance, reputation and reviews, and accessibility.
Example Responses
The example full response generated through visual-only prompting in Figure 1 is as follows:
@@ Appendix A Appendix
The example full response generated through visual-textual prompting in Figure 1 is as follows:
@@
The example full response generated through cross-reflection prompting in Figure 1 is as follows:

### v2

@@ 5.1 Comparison with Baseline Prompting Strategies
MovieLens-1M Amazon-Software Precision@10 Recall@10 NDCG@10 Precision@10 Recall@10 NDCG@10 Text + Image 0.2781 0.2551 0.3523 0.0467 0.4182 0.2915 Rec-GPT4V Liu et al. (2024c) 0.2787 0.2547 0.3542 0.0467 0.4171 0.2964 CoT Wei et al. (2022) 0.2789 0.2540 0.3540 0.0487 0.4354 0.3227 X-Reflect 0.2807 0.2539 0.3577 0.0491 0.4377 0.3291 (+0.93) (-0.47) (+1.53) (+5.14) (+4.66) (+12.90)
5.2 Generalizability of the Proposed Framework
To evaluate the generalizability of our framework, we replace the GPT-4o backbone with another widely used MLLM—LLaVA—while keeping all experimental settings consistent across baselines. Table 3 reports the corresponding recommendation performance. With LLaVA as the backbone, X-Reflect continues to outperform other prompting baselines overall across both datasets. However, all multimodal prompting methods exhibit a decline in Recall@10 on the MovieLens-1M dataset compared to the Text+Image baseline, which directly uses CLIP image embeddings. This suggests that LLaVA may have limited multimodal reasoning capabilities relative to GPT-4o—a result that aligns with expectations.
MovieLens-1M Amazon-Software Precision@10 Recall@10 NDCG@10 Precision@10 Recall@10 NDCG@10 X-Reflect 0.2847 0.2619 0.3621 0.0521 0.4654 0.3266 Visual-textual 0.2803 0.2573 0.3572 0.0490 0.4365 0.3267 (-1.55) (-1.76) (-1.35) (-5.96) (-6.21) (+0.03) Visual-only 0.2844 0.2618 0.3614 0.0478 0.4268 0.3174 (-0.11) (-0.04) (-0.19) (-8.25) (-8.29) (-2.82) Text-only 0.2827 0.2573 0.3604 0.0503 0.4483 0.3371 (-0.70) (-1.76) (-0.47) (-3.45) (-3.67) (+3.21)
5.3 Ablation on Prompts
To better understand the impact of prompt design in multimodal recommendation, we conduct an ablation study comparing our method against three baselines: (1) Text-only prompting (LLM-Rec Lyu et al. (2024)), which generates item descriptions using only textual input; (2) Visual-only prompting (Rec-GPT4V Liu et al. (2024c)), which generates descriptions based solely on the item image; and (3) Visual-textual prompting, which incorporates both the image and the original item description but does not explicitly instruct the model to reason about consistencies or conflicts between the two. The visual-textual baseline uses the following prompt:
“The item description is as follows: {item description}. Please describe this item based on both the description and the input image. Your response should highlight how combining both the text and image provides a comprehensive understanding of the item.”
We perform this experiment on both datasets using GPT-4o as the MLLM backbone. Results show that our method, which explicitly instructs the model to identify and reconcile supportive and conflicting information across modalities, overally achieves the best performance. Surprisingly, the visual-textual prompting baseline performs the worst, even underperforming the unimodal approaches. While counterintuitive, this may be attributed to the generic nature of the generated responses when both modalities are presented without focused guidance. In contrast, unimodal prompts encourage the model to attend more closely to the details within a single modality.
These findings underscore the importance of carefully crafting multimodal prompts to elicit complementary and discriminative information, rather than overwhelming the model with vague or overly broad instructions.
5.4 Complexity Analysis
Let and denote the text input length (in tokens) of the text and image prompts, respectively. Let and represent the tokens generated by the model in response to the first-stage prompts, and let denote the tokens generated in the final stage. For simplicity, we omit the cost of fixed system tokens from the analysis.
We define the latency of a single model call as , where the dominant cost arises from the autoregressive generation output tokens.
Parallel Execution
If the first two model calls (text and image prompts) are executed in parallel, the total latency is approximately:
Sequential Execution
If the first two stages are run sequentially, the total latency becomes:
Note that the final-stage prompt scales with , meaning longer intermediate responses can significantly increase end-to-end latency.
In real-time recommendation systems, minimizing latency is essential. To reduce runtime overhead, we propose limiting the length of the first-stage responses. Specifically, instead of prompting the model to generate full descriptions for the image, we instruct it to produce a concise list of keywords. The prompt is revised as:
“Describe the image using a list of keywords.”
Additionally, to further reduce cost, we substitute the base GPT-4o model with the lighter-weight GPT-4o-mini model for both the image keyword summarization and the final reflection step. All other components remain unchanged.
We refer to this variant as X-Reflect-keyword and evaluate it on the two benchmark datasets. Results are shown in Table 5. Despite reasoning only over keyword-based image summaries, this approach maintains competitive performance compared to the original variant that reasons over full descriptions. More importantly, it significantly reduces the number of input tokens (, one-sided t-test). On average, the word counts of the image descriptions in MovieLens-1M and Amazon-Software are 74.81 and 35.47, respectively, whereas the keyword-based summaries reduce them to 38.68 and 17.52 — nearly a 50% reduction.
MovieLens-1M Amazon-Software Precision@10 Recall@10 NDCG@10 Precision@10 Recall@10 NDCG@10 X-Reflect 0.2847 0.2619 0.3621 0.0521 0.4654 0.3266 X-Reflect-keyword 0.2855 0.2616 0.3635 0.0511 0.4571 0.3393
5.5 Relationship between Text-Image Dissimilarity and Recommendation Performance
To explore how the alignment between an item’s text and image affects recommendation performance, we analyze the text–image dissimilarity and correlate it with per-user recommendation effectiveness. We use the MovieLens-1M dataset as an example.
Computing Text-Image Dissimilarity
We first encode the original item description using a pre-trained sentence encoder, Sentence-BERT Reimers and Gurevych (2019). Separately, we prompt GPT-4o with the visual-only instruction: “Describe the image.” The model generates a natural language description based solely on the visual content. This response is the encoded using the same Sentence-BERT model to ensure consistency in the embedding space.
We compute the cosine dissimilarity (i.e., ) between the two embeddings. One is derived from the item description and the other is derived from the GPT-4o-generated image description. This results in a scalar dissimilarity scrore for each item.
Aggregating User-Level Dissimilarity
For each user, we identify all the items they interacted with in the testing set. We then compute the average text-image dissimilarity across these items, yielding a single dissimilarity score per user. This score captures the overall degree of discrepancy between textual and visual information for the user’s engaged items.
Measuring Recommendation Performance
To evaluate recommendation quality, we compute NDCG@10 for each user under different prompting strategies. This user-level performance metric is the nplotted against the corresponding user-level dissimilarity score.
Observations and Implications
As shown in Fig. 1, the relationship between recommendation performance and text-image dissimilarity exhibits a U-shaped pattern: performance tends to be lower when dissimilarity is either very low (strong disagreement) or very high (strong conflict), and highest when the dissimilarity is moderate.
When the item description and image are highly aligned (i.e., low dissimilarity), both modalities express similar or even redundant information. There is limited benefit from multimodal reasoning, since the image provides little new information beyond what is already stated in the text. The modal may be wasting inference capacity on reconciling modalities that are already in agreement, which adds complexity without improving utility.
On the other end, when dissimilarity is high, the textual and visual inputs are semantically contradictory or incoherent with each other. The model may face difficulty in reconciling conflicting information, leading to uncertainty or hallucinations in the final response. This situation may reflect noisy or misleading data (e.g., misaligned product descriptions and images), where neither modality can be trusted fully.
This insights a practical pathway for inference-time optimization. Instead of applying multimodal prompting indiscriminately, we can follow an adaptive strategy that reserves multimodal prompting for items with moderate dissimilarity, and applies more efficient unimodal methods when the text and image either strongly agree or clearly diverge. Such a strategy can preserve accuracy while reducing inference latency, enabling smarter deployment of multimodal LLMs in recommendation systems.
6 Discussions and Conclusions
In this paper, we introduce X-Reflect, a novel framework that leverages cross-reflection prompting to enhance item representations by jointly integrating textual and visual modalities. Unlike conventional prompting strategies that treat modalities independently, X-Reflect encourages Multimodal Large Language Models to reason across modalities by explicitly identifying, contrasting, and reconciling supportive or conflicting information between text and image.
Our experimental results on two widely used recommendation benchmarks, MovieLens-1M and Amazon-Software, reveal several key findings. Multimodal prompting methods that incorporate both text and image generally outperform text-only prompting methods, highlighting the value of leveraging multiple modalities. Among multimodal methods, X-Reflect consistently outperforms other baselines such as Rec-GPT4V, CoT, and an ablated visual-textual prompting strategy. This improvement is likely to stem from its specific instruction to perform cross-modal reasoning.
We also analyze the relationship between text–image dissimilarity and recommendation performance, uncovering a U-shaped pattern: performance is highest when the two modalities offer moderately dissimilar, complementary information, and lowest when they are either redundant or strongly conflicting. This insight motivates an adaptive inference strategy that applies multimodal reasoning selectively—only when such reasoning is expected to yield the greatest benefit.
To further address practical concerns of latency in real-time systems, we propose a simplified variant called X-Reflect-keyword. Instead of generating full image descriptions, we prompt the model to summarize the image using a concise list of keywords. We find that X-Reflect-keyword achieves comparable performance while significantly reducing input length—cutting the average number of words in image descriptions by nearly 50 percent. This offers a promising path toward cost-effective deployment without compromising recommendation quality.
Overall, X-Reflect demonstrates that explicit multimodal reflection is a powerful mechanism for bridging visual and textual understanding in recommendation tasks. Future work will explore extensions of this framework to user-centric modeling and investigate the use of cross-reflection prompting with additional modalities such as audio or user interaction context.
@@ 6 Discussions and Conclusions
However, the proposed framework is not without its limitations. One notable drawback is the increased computational overhead associated with the need for multiple MLLM inferences per item. This requirement can be resource-intensive, especially when processing large-scale datasets or deploying the system in real-time applications. To mitigate this issue, we introduce a simplified variant of our framework. We also show that future research could explore optimization strategies, such as selectively applying cross-reflection prompting to a subset of items that are likely to benefit the most from this process. Additionally, advances in model efficiency, such as improved inference algorithms or more lightweight MLLM architectures, could further reduce the computational burden.
Appendix A Example Responses
The example full response generated through visual-only prompting in Fig. 2 is as follows:
@@ Appendix A Example Responses
The example full response generated through visual-textual prompting in Fig. 2 is as follows:
@@
The example full response generated through cross-reflection prompting in Fig. 2 is as follows:
@@ Conclusion:
Appendix B KAR Augmentation Details
In the work by Xi et al. (2023), a specific prompt was designed to elicit factual knowledge about items. The prompt instructed the model to:
“Introduce {item}, {item description} and describe its attributes precisely (including but not limited to {scenario-specific factors})”. The {scenario-specific factors} were identified by prompting an LLM with the following instruction: “List the importance factors or features that determine whether a user will be interested in a {item}.”
For the MovieLens-1M dataset, we directly use the factual knowledge about movies generated in their study. For the Amazon-Software dataset, we apply the same approach to first identify the {scenario-specific factors}. The prompt was adapted to:
“List the importance factors or features that determine whether a user will be interested in software.”
The set of factors generated by ChatGPT includes functionality, user interface, user experience, performance, compatibility, cost, security, support and documentation, customizability, scalability, reliability, updates and maintenance, reputation and reviews, and accessibility.